Assume that you contribute $340 per month to a retirement plan for 25 years. Then you are able to increase the contribution to $540 per month for the next 25 years. Given a 7.2 percent interest rate, what is the value of your retirement plan after the 50 years? (Do not round intermediate calculations and round your final answer to 2 decimal places.)

Find Future value of multiple annuities

Please answer correctly
Answer #1

We can use the formula of future value of annuity (FVA) to find out the amount after 50 years.

Formula:

|(1 + i/a)na - 1 FVA = A i/a

Where,

A = Monthly contribution

i = rate of interest

n= number of periods

a= number of months in a year

So since it is a question of multiple annuities, when we put the values in the formula we will get:

FVA = 340 (1 + 0.072/12,5(! - 17540||(1 +0.072/22730.1! - 11 - [(1 +0.072/12)41! _ 1]] 0.072/12 0.072/12 0.072/12

= 340(836.1995) + 540(5867.7762-836.1995)

- 284307,82 +2717051.45

- 300139.27

That means after 50 years your retirement account will show a balance of $3,001,359,27

Note: I have not rounded off intermediate calculations in my calculations. I have only rounded off the final answer to 2 decimals.